OpenStudy (anonymous):

write the equation of the line in slope-intercept form that is parallel to the given line and passes through the given point. y=4/9x+1, (45,18)

OpenStudy (atlas):

The slope-intercept form of line is y=mx +c as I mentioned earlier If a line is parallel to another line their slopes are same.

OpenStudy (atlas):

You can find the slope of the given line : m by matching it with the standard form of slope intercept y = mx +c y = (4/9)x +1

OpenStudy (atlas):

can you tell me the slope of the given line

OpenStudy (anonymous):

4/9

OpenStudy (atlas):

right!!

OpenStudy (atlas):

so the new line which is PARALLEL to the given line will have the same slope: SO, the equation of the new line is y =4/9 x +c we got m - now we just need to get c

OpenStudy (atlas):

If the line is passing through (45,18) the points must satisfy the equation for the line we will put x = 45 and y=18 in the above equation and find c

OpenStudy (anonymous):

so it would be y=4/9x+2 ?

OpenStudy (anonymous):

@tgawade is this the right answer y=4/9x+2 ?